If 6 men can do a piece of work in 14 days, how many men are needed to do the work in 21 days?
the Answer was given as 4. can you show me how this answer was reached?

Answer by John10(297)

If 6 men can do a piece of work in 14 days, how many men are needed to do the work in 21 days?
-------------------------------
As you know that the larger group of people will complete faster than smaller group of people. Thus it is relatively inverse
Let D is number of days and N is number of men
N = k/D
6 = k/14
==>k = 6 * 14 = 84
If we have D = 21 days then
N = k/D = 84/21 = 4 men.
Hope it helps! John